OpenStudy (anonymous):

if a^2+b^2=7ab . Prove that : 2ln((a+b)/3)=lna+lnb.

OpenStudy (zarkon):

\[2\ln\left(\frac{a+b}{3}\right)\] \[=\ln\left(\left[\frac{a+b}{3}\right]^2\right)\] \[=\ln\left(\frac{a^2+b^2+2ab}{3^2}\right)\] \[=\ln\left(\frac{7ab+2ab}{9}\right)\] \[=\ln\left(\frac{9ab}{9}\right)\] \[=\ln\left(ab\right)\] \[=\ln(a)+\ln(b)\]